solve the system by the elimination method
5x + 5y = -11
7x - 3y = 17
The solution is?
Please show work if possible - I came up with 2 different answers for y
-81/25 or -162/31
Thank you for your help

Answer by stanbon(75887)

solve the system by the elimination method
5x + 5y = -11
7x - 3y = 17
-------------------------
35x + 35y = -77
35x - 15y = 85
----------------------
50y = -162
y = -81/25
-------------------
Substitute and solve for "x":
5x + 5(-81/25) = -11
5x - 81/5 = -11
5x = 26/5
x = 26/25
===================
Cheers,
Stan H.